If the current total inventory value across 10 warehouses is $6500000, and the company plans to expand it's operations to 13 warehouses, what would be your best estimate of the new total inventory value?
Answer #1

Number of existing warehouses(n1) = 10

Number of future warehouses (n2) = 13

Current total inventory (x1) = $6500000

New total inventory value = x1 × √(n2/n1)

= 6500000 / √(13/10)

= 6500000 / √1.3

= 6500000/1.140175425

= $5700877